# IGNN-Solver: A Graph Neural Solver for Implicit Graph Neural Networks

### v1

@@ 4.2 Training the IGNN-Solver
Initializer Loss Functions.
@@
Reconstruction Loss Functions.
@@
Auxiliary Loss Functions.
@@
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ control the weights of three loss functions mentioned above and elaborate more in Appendix C.3.
5 Experiments
In this section, we compare the speed/accuracy Pareto curve rather than a single point on the curve of IGNN-Solver with IGNNs and several state-of-the-art (SOTA) GNNs on various graph classification tasks at both node and graph levels. We aim to show that 1) IGNN-Solver can achieve nearly a 1.5×1.5$\times$1.5 × to 8×8$\times$8 × inference acceleration without sacrificing accuracy, and 2) IGNN-Solver are extremely compact and add little overhead to training.
Specifically, we compare our approach against $12$ representative baselines and $2$ variants: IGNN w. AA (using original AA to accelerate IGNN), and IGNN w. NN (using a standard NN solver to replace our proposed graph neural solver for parameter learning) on $9$ different-field and real-world node classification tasks, including $4$ citation datasets: Citeseer, ACM, CoraFull, ogbn-arxiv, $3$ social interaction datasets: BlogCatalog, Flickr, Reddit, and $2$ product network datasets Amazon-all, ogbn-products. More details about the datasets are given in Appendix C.1. Notably, to demonstrate the scalability of the proposed model to larger datasets, we use $4$ large-scale datasets: Amazon-all, Reddit, ogbn-arxiv and ogbn-products, two of which are from the Open Graph Benchmark (OGB) [22] [^1] The code to reproduce the results in this section has been uploaded as supplementary material..
Additionally, in Section 5.2, we demonstrate that the training overhead on IGNN-Solver is minimized relative to the total time overhead on IGNN. In Section 5.3, we provide additional evidence regarding the convergence and generalizability of the IGNN-Solver. In Section 5.4, we conduct an ablation study to validate the stability of the IGNN-Solver and the effectiveness of its components. Experimental setups, descriptions and additional results are detailed in Appendices BC and D.
@@ 5 Experiments
Figure 2: Comparison of inference time with IGNN-Solvers on $4$ large-scale datasets. More results on Citeseer, ACM, CoraFull, BlogCatalog, and Flickr can be found in Figure 6, Appendix D.1. All speed/accuracy curves within the plot are benchmarked on the same GPU with the same experimental setting, averaged over $5$ independent runs.
Table 1: Node classification accuracy results on $4$ large-scale real-world datasets (more results on small-scale datasets including Citeseer, ACM, CoraFull, BlogCatalog, Flickr can be found in Table 3, Appendix D.1), with experiments conducted over five trials. The mean accuracy ± standard deviation (%) is reported. The best and the runner-up results are highlighted in boldface and underline, respectively.
@@
IGNN w. AA $83.44\pm 0.19$ $92.37\pm 0.35$ $71.73\pm 0.41$ $74.61\pm 0.28$
IGNN w. NN $84.13\pm 1.04$ $\underline{92.42}\pm 0.41$ $70.78\pm 0.37$ $\underline{74.69}\pm 0.31$
IGNN-Solver $\textbf{84.50}\pm 0.70$ $\textbf{93.91}\pm 0.31$ $\textbf{72.53}\pm 0.41$ $\textbf{74.90}\pm 0.20$
5.1 Performance and Efficiency Comparison
In order to demonstrate the superiority of the IGNN-Solver over IGNNs in terms of both performance and efficiency, we specifically analyze the movement of the speed/accuracy Pareto curve across various datasets rather than concentrating on a single point on the curve, which is depicted in Figure 2 and 6 (see Appendix D.1). All experiments are conducted five times, and the best results are reported. The training procedure details and hyperparameters used in each task are provided in Appendix C.
From Figure 2 (results on large-scale datasets: Amazon-all, Reddit, ogbn-arxiv and ogbn-products), we can observe that: 1) Regarding comprehensive performance and efficiency, IGNN-Solver generally outperforms all other methods, especially in large datasets with a greater graph radius, where this advantage is often more pronounced. This is attributed to the IGNN-Solver’s significant improvement in the convergence path and the enhancement of the fixed-point equation’s initial point through the initializer; 2) The rapid inference advantage of the IGNN-Solver. For instance, in the large-scale dataset Amazon-all, our IGNN-Solver achieves 8×8$\times$8 × faster inference speed than IGNN while maintaining the same performance, and there is consistently at least 1.5×1.5$\times$1.5 × acceleration, and a similar pattern has been observed as well on other datasets. Please see more results on other small-scale datasets in Figure 6 of Appendix D.1.
Furthermore, we present the node classification performance results for each method on large-scale datasets(including Amazon-all, Reddit, ogbn-arxiv and ogbn-products) in Table 1 and other small-scale datasets in Table 3 (see Appendix D.1). All experiments are conducted five times, and the average results along with the standard deviation are reported. The training procedure details and hyperparameters used in each task are provided in Appendix C. It can be observed that among all methods, implicit GNNs with infinite depth are often superior to shallow explicit GNNs in most cases. Besides, IGNN-Solver consistently shows higher accuracy percentages across most datasets compared to other SOTA explicit GNNs like DEMO-Net [48], GCNII [11], ACM-GCN [35] etc, which indicates their superior performance.
5.2 Model Size in Training
(a) Overhead proportion on Citeseer.
(b) Overhead proportion on Amazon-all.
Figure 3: Relative parameter size and training time of IGNN-Solver on small dataset Citeseer and large one Amazon-all, with similar patterns observed in other datasets.
To investigate the proportion of training overhead attributed to the solver throughout the entire training process, we depict in Figure 3 the percentage of training computations on IGNN-Solver relative to the total time overhead on IGNN. Our approach is not only effective but also incurs minimal training overhead for the solver: the solver module is relatively small, and requires only about $1\%$ of the original training time required by the IGNN model. Besides, this proportion will be lower for large-scale data. For instance, on the Amazon dataset, IGNN necessitates a total runtime of $3$ hours, whereas the solver only requires approximately $1.6$ minutes, indicating that our solver maintains its lightweight nature across any-scale datasets.
5.3 More Efficiency Study
Figure 4: The convergence curve of IGNN-Solver and IGNN on ogbn-arxiv. The former improves the convergence rate.
We provide additional evidence on the convergence and generalizability of the neural solvers in Figure 4, where we compare the convergence of a pre-trained IGNN model under 1) canonical iteration in IGNN [20]; and 2) IGNN-Solver with $K=6$.
From Figure 4, we first note that the IGNN-Solver trained with $K$ unrolling steps can generalize beyond $K$, and both solvers eventually reach a stable state, demonstrating good convergence. Secondly, thanks to the fixed-point solution of the graph neural parameterization, we observe that our IGNN-Solver continuously enhances the convergence of the canonical iterators while being more computationally efficient (i.e., each step of the neural solver is cheaper than the standard iterator step). This explains the observed improvement in inference efficiency of at least 1.5×1.5$\times$1.5 × in Section 5.1.
Figure 5: Ablation studies for IGNN-Solver.
5.4 Ablation Study
Here, we analyze the effect of different loss components of the IGNN-Solver. For the main convergence loss $\mathcal{L}_{\text{conv}}$ of fixed-point iterations in equation 11, we designed two contrasting schemes concerning its weight $\lambda_{1}^{[k]}$s: setting them to a constant value (i.e., $\lambda_{1}^{[k]}=\lambda_{1}$ for all $k$), or setting all values except the $K$-th term to zero (i.e., $\lambda_{1}^{[k]}=\lambda_{1}$ if $k=K$ else 00). The results of IGNN-Solver and its variants on Citeseer are shown in Figure 5. The patterns observed in other datasets and solvers exhibit similarities.
It is observed that the two variant solvers still perform well, yet our suggested monotonically increasing scheme (i.e., emphasizing the later steps to a greater extent) exhibits the best performance. Additionally, removing the initializer or the alpha loss $\mathcal{L}_{\alpha}$ will correspondingly affect performance, with the former having a more detrimental impact on the solver. Nonetheless, the performance of all these ablation settings remains significantly superior to methods without a solver, indicating the advantages of employing a custom learnable solver for the IGNN model.
6 Conclusion
This paper introduces IGNN-Solver, a novel graph neural solver tailored for achieving fast fixed-point solving in IGNNs. By leveraging the generalized Anderson Acceleration method and parameterizing it with a tiny GNN, IGNN-Solver learns iterative updates as a graph-dependent temporal process. Our extensive experiments demonstrate the significant acceleration in inference achieved by IGNN-Solvers, with a speedup ranging from 1.5×1.5$\times$1.5 × to 8×8$\times$8 ×, all while maintaining accuracy. Notably, this acceleration is particularly pronounced as the scale of the graph increases. These findings underscore the potential of IGNN-Solver for large-scale end-to-end deployment in real-world applications.
Limitations.
In contrast to previous works on IGNNs [20; 33; 7], the theoretical guarantee of fixed-point existence and stability in IGNN-Solver remains uncertain, notwithstanding favorable empirical findings in Figure D.2. This is because we make no specific constraint on the formulation of IGNNs, a departure from the approach adopted in previous IGNN-related studies.
Acknowledgments
R. C. Qiu would like to acknowledge the National Natural Science Foundation of China (via NSFC-12141107), the Interdisciplinary Research Program of HUST (2023JCYJ012), the Key Research and Development Program of Guangxi (GuiKe-AB21196034). F. Zhou was supported by the National Natural Science Foundation of China Project (via NSFC-62106121) and the MOE Project of Key Research Institute of Humanities and Social Sciences (22JJD110001).
Z. Ling is supported by the National Natural Science Foundation of China (via NSFC-62406119), the Natural Science Foundation of Hubei Province (2024AFB074), and the Guangdong Provincial Key Laboratory of Mathematical Foundations for Artificial Intelligence (2023B1212010001).
Supplementary Material
Appendix A Anderson Acceleration Solver for Deep Equilibriums (DEQs)
Algorithm 2 Anderson Acceleration Solver
3:for $k=0$, $\ldots$, $K-1$ do
6:     3) ${\bm{Z}}^{[k+1]}=\beta\sum_{i=0}^{m_{k}}\alpha_{i}^{[k]}f_{\theta}({\bm{Z}}^{[k-m_{k}+i]})+(1-\beta)\sum_{i=0}^{m_{k}}\alpha_{i}^{[k]}{\bm{Z}}^{[k-m_{k}+i]}\quad$triangleright AA update
7:end for
Given an implicit network layer $f_{\theta}$ and an input ${\bm{X}}$, DEQs aim to find the fixed point ${\bm{Z}}^{*}$ in the system that represents the equilibrium state of the network output by solving a root-finding problem:
$g_{\theta}\left({\bm{Z}}^{\star},{\bm{X}}\right):=f_{\theta}\left({\bm{Z}}^{\star},{\bm{X}}\right)-{\bm{Z}}^{\star}=0.$ (12)
In practical solving of DEQs, it is often necessary to utilize fixed-point solvers during the training process, such as the classical Broyden method [9] or Anderson Acceleration (AA) [2], which can directly seek the equilibrium point ${\bm{Z}}^{\star}$ through quasi-Newton methods. This approach demonstrates super-linear convergence properties.
We briefly introduce AA-solver here, as our approach relies on it. Algorithm 2 is a pseudocode example illustrating it, where ${\bm{G}}^{[k]}=[g_{\theta}(\mathbf{z}^{[k-m_{k}]})\ldots g_{\theta}(\mathbf{z}^{[k]})]$ is the past residuals. It can be observed that the basic idea of Anderson Acceleration (AA) is to accelerate the generation of the next approximate solution by utilizing the information stored from the history $m$ iteration steps, constructing a normalized linear combination with weights $\alpha^{[k]}$. It computes the weights greedily to minimize the linear combination of past AA update steps. This approach is typically particularly effective for iteratively slow-converging processes, such as solving implicit functions.
Subsequently, during the backward process, the implicit function theorem [28] can be employed to implicitly differentiate through the equilibrium point and generate gradients with respect to the model parameters $\theta$ by solving linear equations based on the Jacobian matrix. Since the computation process relies only on the final output and does not require storing any intermediate activation values [4], the memory consumption during the training process remains constant, which is also a key reason why DEQs are attractive.
Appendix B Training Strategies for IGNN using IGNN-Solver.
B.1 IGNN Training with frozen IGNN-Solver
During the training processing of the IGNN model, we assume that both the initializer $h_{\phi}$ and the neural solver $s_{\xi}$ of IGNN-Solver have approached the desired state via the training method outlined in Section 4.1, and they will no longer be updated in the subsequent model training. For a given IGNN layer $f_{\theta}$, as well as the input graph ${\bm{A}}$ and feature matrix ${\bm{X}}$, we define:
$g_{\theta}({\bm{Z}}^{\star},\hat{{\bm{A}}},{\bm{X}}):=f_{\theta}({\bm{Z}}^{\star},\hat{{\bm{A}}},{\bm{X}})-{\bm{Z}}^{\star}=0,$ (13)
and utilize the IGNN-Solver proposed in Section 4.1 to solve equation 13, obtaining the fixed point ${\bm{Z}}^{\star}$. Furthermore, the fixed point ${\bm{Z}}^{\star}$ is passed through a linear layer to obtain an embedding ${\bm{H}}=\text{Linear}({\bm{Z}}^{\star})$. Then we jointly with downstream tasks, compute the Binary Cross Entropy (BCE) loss [15] to back-propagate and update the parameters of $f_{\theta}$ in a supervised way. The complete process is illustrated in Algorithm 3.
Algorithm 3 IGNN model training
1:graph matrix $\hat{{\bm{A}}}$, feature matrix ${\bm{X}}$, fixed IGNN-Solver $s_{\xi}$ and initializer $h_{\phi}$
3:while Stopping condition is not met do
4:     ${\bm{Z}}^{\star}\leftarrow\text{Solve}\ f_{\theta}({\bm{Z}}^{\star},\hat{{\bm{A}}},{\bm{X}})-{\bm{Z}}^{\star}=0$triangleright via frozen $s_{\xi}$, $\hat{{\bm{A}}}$ and initial value ${\bm{Z}}^{[0]}$
5:     ${\bm{H}}\leftarrow\text{Linear}({\bm{Z}}^{\star})$
7:     Back-propagate $\mathcal{L}_{\text{task}}$ to update $f_{\theta}$
8:end while
9:return $f_{\theta}$
B.2 Advanced IGNN Training Strategy with IGNN-Solver
  1. (i)
  2. (ii)
  3. (iii)
  4. (iv)
Here $T_{1}$ is approximately 4% to 10% of $T_{2}$, adjusted as needed. The sum of all $T_{2}$ values is referred to as $\text{epoch}_{\text{max}}$. It is reassuring that the additional cost of fine-tuning the IGNN-Solver (step (iii)) is sufficiently offset by the substantial benefits of accelerated solving (step (ii)). Moreover, we are surprised to find that the proposed IGNN-Solver does not exhibit a decline in expressive capability for excessively small or large $T_{1}$ values. This indicates that, thanks to its robust stability, accidentally setting $T_{1}$ high does not significantly affect the normal training of the IGNN. Conversely, if $T_{1}$ is set slightly lower, the model still demonstrates good generalization capabilities. Thus, the IGNN-Solver shows low sensitivity to $T_{1}$.
Appendix C Experimental Details
The experiments are conducted on $9$ public datasets of different scales, including $5$ widely adopted benchmark datasets Citeseer, ACM, CoraFull, BlogCatalog, Flickr, and $4$ large-scale benchmark datasets Amazon-all, Reddit, ogbn-arxiv and ogbn-products. We endeavor to conduct inference testing under almost identical hyperparameter conditions as previous work [7; 33; 20; 45; 26; 42; 47; 17; 50], including performance and efficiency comparisons. All the Experiments are conducted independently five times (i.e., using $5$ random seeds) on a machine with Intel(R) Xeon(R) Gold 6138 CPU @ 2.00GHz with a single 3090 GPU. A detailed description of all tasks and additional details are provided below.
C.1 Datasets
  • ACM [54]: It is a citation network dataset, where nodes represent papers and node features are constructed by the keywords. Papers are divided into 3 categories according to their types of conferences.
  • Flickr [48]: It is a graphic social network where nodes represent users and edges correspond to the friendships among users. All the nodes are divided into 9 classes according to the interest groups of users.
Table 2: Hyper-paramaters setting on training. *In addition, we decay the loss weight $\lambda_{1}$ from 00 to the setting value on a linear schedule over all IGNN-Solver training steps.
Small Citeseer 128 0.5 0.002 10 100 360 1000 0.1 5 1e-4
ACM 128 0.5 0.001 10 100 180 1000 0.3 5 1e-4
CoraFull 512 0.5 0.002 15 300 4200 1000 0.3 5 1e-4
BlogCatalog 512 0.5 0.001 10 100 360 1000 0.5 5 1e-4
Flickr 512 0.5 0.002 15 200 540 1000 0.1 5 1e-5
Large Amazon-all 128 0.5 0.005 20 1000 16970 28285 0.3 5 1e-5
Reddit 512 0.5 0.005 15 2000 139,779 46,593 0.5 5 1e-5
ogbn-arxiv 128 0.5 0.001 20 2000 90,941 48,603 0.5 5 1e-5
ogbn-pruducts 128 0.5 0.001 20 2000 196,615 2,213,091 0.5 5 1e-5
C.2 Baselines
We compare IGNN-Solver with $12$ state-of-the-art methods, including $3$ implicit IGNNs, i.e., MIGNN [7], EIGNN [33], IGNN [20], and $9$ explicit/traditional GNNs, i.e., AM-GCN [45], GCN [26], GAT [42], SGC [47], APPNP [17], JKNet [50], DEMO-Net [48], GCNII [11], ACM-GCN [35]. The specific parameter settings are as follows:
For the configuration of hidden layers, for the sake of fairness, we set the same hidden layers for all baselines. For example, in the Flickr dataset, the dimension of the hidden layers is uniformly set to $512$ for all methods (including ours, see in Table 2). In addition, for EIGNN, the arbitrary gamma is set to $0.8$, which is consistent with the original paper. For AM-GCN, the number of nearest neighbors in KNN graph is set from $5,6,7$. For GCN, the hidden layer setup is same as others. For GAT, three attention headers are used for each layer. For SGC, the power of self-loops in the graph adjacency matrix is set to $3$. For APPNP, we set the number of iterations to $3$ and teleport probability $\alpha$ to $0.5$. For JKNet, we set layers to $1$ in small datasets and set to $8$ in large ones. For DEMO-Net, the regularization parameter is set as $45e-$ 4, the learning rate is set as $35e-$ 3, and the hash dimension is set as 256. For GCNII, we set $\alpha_{\ell}=0.1$ and $L_{2}$ regularization to $45e-$ 4, consistent with the original paper. For ACM-GCN, we set layers to $1$ in small datasets and set to $4$ in large ones. For our IGNN-Solver and its variants, we employ the same settings as the basic IGNN model. Furthermore, we adjust these parameters affecting the convergence and performance of IGNN through a combined approach of hierarchical grid search and manual tuning. The Adam optimizer [25] is used for optimization.
C.3 Hyper-parameters Setting in IGNN-Solver
We present in Table 2 all the hyper-parameters preset by IGNN-Solver across all datasets, where $K$ represents the threshold of maximum iteration in IGNN-Solver, $\text{epoch}_{\text{max}}$ means the maximum training steps in IGNN. It is worth noting that we set the value of $\lambda_{1}$ in a linearly increasing manner to penalize the intermediate estimation error when the training step $K$ is large. Similarly, we set the value of $\lambda_{3}$ in a linearly decreasing manner to prevent overfitting $\alpha$ during later training epochs.
@@ C.3 Hyper-parameters Setting in IGNN-Solver
Figure 6: Comparison of inference time with IGNN-Solvers on $5$ small-scale datasets. All speed/accuracy curves within the same plot are benchmarked on the same GPU with the same experimental setting, averaged over $5$ independent runs.
Table 3: Node Classification accuracy results on $5$ small-scale real-world datasets, with experiments conducted over five trials. The mean accuracy ± standard deviation (%) is reported. The best and the runner-up results are highlighted in boldface and underline, respectively.
@@
IGNN w. AA $\underline{75.28}\pm 0.38$ $\textbf{91.34}\pm 0.46$ $\underline{65.88}\pm 0.34$ $76.82\pm 0.34$ $75.84\pm 0.27$
IGNN w. NN $74.78\pm 0.42$ $91.08\pm 0.56$ $65.62\pm 0.16$ $76.88\pm 0.74$ $78.55\pm 0.49$
IGNN-Solver $\textbf{75.60}\pm 0.22$ $\underline{91.20}\pm 1.15$ $\textbf{66.08}\pm 1.23$ $\underline{77.64}\pm 0.54$ $\textbf{80.14}\pm 0.23$
Appendix D More Results and Additional Experiments
D.1 More Performance and Efficiency Comparison on other small-scale tasks

### v2

@@ IV-C Training IGNN-Solver
Initializer Loss Functions
@@
Reconstruction Loss Functions
@@
Auxiliary Loss Functions
@@
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ control the weights of three loss functions mentioned above and elaborate more in Section A-B.
@@
Figure 2: Speed-accuracy tradeoff curves comparing the inference time of IGNN with IGNN-Solver and its variants IGNN-Solver (NN) and IGNN-Solver (AA), across four large-scale datasets: Amazon-all (2(a)), Reddit (2(b)), ogbn-arxiv (2(c)) and ogbn-products (2(d)). Each plot shows the average performance over five independent runs under identical experimental conditions.
TABLE I: Node classification accuracy results on four large-scale real-world datasets (more results on small-scale datasets including Citeseer, ACM, CoraFull, BlogCatalog, Flickr can be found in Table II), with experiments conducted over five trials. The mean accuracy (%) ± standard deviation is reported. The best and the runner-up results are highlighted in boldface and underline, respectively.
@@
IGNN-Solver (AA) $83.44\pm 0.19$ $92.37\pm 0.35$ $71.73\pm 0.41$ $74.61\pm 0.28$
IGNN-Solver (NN) $84.13\pm 1.04$ $\underline{92.42}\pm 0.41$ $70.78\pm 0.37$ $\underline{74.69}\pm 0.31$
IGNN-Solver (Ours) $\textbf{84.50}\pm 0.70$ $\textbf{93.91}\pm 0.31$ $\textbf{72.53}\pm 0.41$ $\textbf{74.90}\pm 0.20$
V Experiments
In this section, we compare the speed/accuracy Pareto curve rather than a single point on the curve of IGNN-Solver with IGNNs and several state-of-the-art (SOTA) GNNs across various graph classification tasks, both at the node and graph levels. Our goal is to demonstrate: 1) IGNN-Solver achieves nearly a 1.5×1.5$\times$1.5 × to 8×8$\times$8 × inference acceleration without sacrificing accuracy; and 2) IGNN-Solver is extremely compact, adding little overhead to training. Specifically, we compare our approach with numerous representative baselines and two variants of our method: IGNN w. AA (using the original Anderson Acceleration to speed up IGNN) and IGNN w. NN (replacing our proposed graph neural solver with a standard neural network solver for parameter learning).
In Section V-D, we sequentially demonstrate that the training overhead on IGNN-Solver is little relative to the total time overhead on IGNN. Additionally, we provide further evidence regarding the convergence and generalizability of the IGNN-Solver. We compare the training/inference time costs per epoch between IGNN and IGNN-Solver, visualizing the advantages of our approach. Furthermore, we conduct ablation studies to validate the stability of the IGNN solver and assess the effectiveness of its components. Additional experimental setups and descriptions are detailed in Appendix A.
@@ V Experiments
Figure 3: Speed-accuracy tradeoff curves comparing the inference time of IGNN with IGNN-Solver and its variants IGNN-Solver (NN) and IGNN-Solver (AA), across five small-scale datasets: Citeseer (3(a)), ACM (3(b)), CoraFull (3(c)), BlogCatalog (3(d)) and Flickr (3(e)). Each plot represents the average performance over five independent runs under identical experimental conditions.
TABLE II: Node Classification accuracy results on five small-scale real-world datasets, with experiments conducted over five trials. The mean accuracy (%) ± standard deviation is reported. The best and the runner-up results are highlighted in boldface and underline, respectively.
@@
IGNN-Solver (AA) $\underline{75.28}\pm 0.38$ $\textbf{91.34}\pm 0.46$ $\underline{65.88}\pm 0.34$ $76.82\pm 0.34$ $75.84\pm 0.27$
IGNN-Solver (NN) $74.78\pm 0.42$ $91.08\pm 0.56$ $65.62\pm 0.16$ $76.88\pm 0.74$ $78.55\pm 0.49$
IGNN-Solver (Ours) $\textbf{75.60}\pm 0.22$ $\underline{91.20}\pm 1.15$ $\textbf{66.08}\pm 1.23$ $\underline{77.64}\pm 0.54$ $\textbf{80.14}\pm 0.23$
V-A Experiments Settings
V-A1 Datasets
The evaluation is conducted on nine different-field, real-world benchmarks for node classification tasks, including four citation datasets (Citeseer, ACM, CoraFull, ogbn-arxiv), three social interaction datasets (BlogCatalog, Flickr, Reddit), and two product network datasets (Amazon-all, ogbn-products); as well as five bioinformatics benchmarks (MUTAG, PTC_MR, COX2, PROTEINS, NCI1) for graph classification tasks. Further details can be found in Appendix A. Notably, to demonstrate the scalability of the proposed model on larger datasets, we use four large-scale datasets: Amazon-all, Reddit, ogbn-arxiv, and ogbn-products, two of which are from the Open Graph Benchmark (OGB) [22].
V-A2 Baselines
We compare IGNN-Solver with a comprehensive set of baselines in graph node classification tasks, including three implicit GNNs, i.e., MIGNN [3], EIGNN [2], IGNN [1], and nine explicit/traditional GNNs, i.e., AM-GCN [25], GCN [12], GAT [23], SGC [48], APPNP [49], JKNet [50], DEMO-Net [51], GCNII [52], ACM-GCN [53]. For the sake of fairness, we use the same hidden layer configuration for all baseline models. For example, in the Flickr dataset, the dimension of the hidden layers is uniformly set to $512$ for all methods (including ours, see in Table IV). In addition, for EIGNN, the arbitrary gamma is set to $0.8$, which is consistent with the original paper. For AM-GCN, the number of nearest neighbors in the KNN graph is set from $5,6,7$. For GCN, the hidden layer setup is the same as that for others. For GAT, three attention headers are used for each layer. For SGC, the power of self-loops in the graph adjacency matrix is set to $3$. For APPNP, we set the number of iterations to $3$ and teleport probability $\alpha$ to $0.5$. For JKNet, we set layers to $1$ in small datasets and set to $8$ in large ones. For DEMO-Net, the regularization parameter is set as $45e-$ 4, the learning rate is set as $35e-$ 3, and the hash dimension is set as $256$. For GCNII, we set $\alpha_{\ell}=0.1$ and $L_{2}$ regularization to $45e-$ 4, consistent with the original paper. For ACM-GCN, we set layers to $1$ in small datasets and set to $4$ in large ones. For our IGNN-Solver and its variants, we employ the same settings as the basic IGNN model. Furthermore, we adjust these parameters affecting the convergence and performance of IGNN through a combined approach of hierarchical grid search and manual tuning. The Adam optimizer [54] is used for optimization.
V-B Performance on Node Classification Tasks
V-B1 Experimental setup
To demonstrate the superiority of the IGNN-Solver over IGNNs in terms of both performance and efficiency, we analyze the movement of the speed/accuracy Pareto curve across various datasets, rather than concentrating on a single point on the curve, which is depicted in Figures  2 and 3. All experiments are conducted five times, and the best results are reported. Furthermore, we present the node classification performance results for each method on various datasets, which are shown in Tables I and II. All experiments are conducted five times, and we report the average results along with the standard deviation. Additional details on the training procedure and the hyperparameters used for each task are provided in Appendices A and B.
V-B2 Results on large-scale tasks
From Figure 2 (the speed/accuracy Pareto curve on large-scale datasets: Amazon-all, Reddit, ogbn-arxiv and ogbn-products), we observe that:
  1. 1.
    Comprehensive performance and efficiency: IGNN-Solver generally outperforms all other methods, particularly in large datasets with a greater graph radius. This advantage is more pronounced in such cases. The improved performance is attributed to IGNN-Solver’s significant enhancement in the convergence path and the better initialization of the fixed-point equation through the initializer;
  2. 2.
    Rapid inference advantage: IGNN-Solver demonstrates a notable speed advantage during inference. For instance, on the large-scale dataset Amazon-all, our IGNN-Solver achieves 8×8$\times$8 × faster inference speed than IGNN while maintaining the same performance, and there is consistently at least 1.5×1.5$\times$1.5 × acceleration and a similar pattern has been observed as well on other datasets.
From Table I, we observe that implicit GNNs with infinite depth generally outperform shallow explicit GNNs in most cases. Furthermore, IGNN-Solver consistently shows higher accuracy percentages across most datasets compared to other state-of-the-art explicit GNNs like DEMO-Net [51], GCNII [52] etc, indicating its superior performance.
V-B3 Results on small-scale tasks